OpenStudy (anonymous):

For which system of equation is (5,3) the solution?

mathslover (mathslover):

There must be some options?

OpenStudy (anonymous):

Hold on.

mathslover (mathslover):

No problem.

OpenStudy (anonymous):

A. 3x-2y=9 3x+2y=14 B. x-y=-2 4x-3y=11 C.-2x-y=-13 x+2y=-11 D. 2x-y=7 2x+7=31

mathslover (mathslover):

In each equation , put x = 5 and y = 3, if both are satisfied then you will get the answer.

mathslover (mathslover):

For example : A.) 3x - 2y = 9 3x + 2y = 14 => 3(5) - 2(3) = 9 3(5) + 2(3) = 14 => 9 = 9 (satisfied) 21 = 14 (NOT satisfied) So , A) is not ur choice. Do the same for all...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):